OpenStudy (anonymous):

The sum of 4 times a number and 14 is 16. What is the number?

jimthompson5910 (jim_thompson5910):

The sum of 4 times a number and 14 is 16 translates to 4x + 14 = 16

jimthompson5910 (jim_thompson5910):

does that help?

OpenStudy (anonymous):

\[-10\]

OpenStudy (anonymous):

not really, 4 times a number+ 14 cat be 16, unless its a negative

OpenStudy (anonymous):

@Giant_Cupcake the answer is \[-10\]

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

Nvm, x = \[\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

4 x .5 = 2 + 14 = 16 Please give me medal :3

OpenStudy (anonymous):

Yeah your right because multiplying by a half is the same as dividing How do you give someone a medal?

OpenStudy (anonymous):

Nvm i figures it out

OpenStudy (anonymous):

Looks like you found it. :p

OpenStudy (anonymous):

i did lol